# Adapting Pre-Trained Vision Models for Novel Instance Detection and Segmentation

### v1

@@ Abstract
Novel Instance Detection and Segmentation (NIDS) aims at detecting and segmenting novel object instances given a few examples of each instance. We propose a unified framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment. Leveraging recent advancements in large vision methods, we utilize the Grounding DINO and Segment Anything Model (SAM) to obtain object proposals with accurate bounding boxes and masks. Central to our approach is the generation of high-quality instance embeddings. We utilize foreground feature averages of patch embeddings from the DINOv2 ViT backbone, followed by refinement through a weight adapter mechanism that we introduce. We show experimentally that our weight adapter can adjust the embeddings locally within their feature space and effectively limit overfitting. This methodology enables a straightforward matching strategy, resulting in significant performance gains. Our framework surpasses current state-of-the-art methods, demonstrating notable improvements of 22.3, 46.2, 10.3, and 24.0 in average precision (AP) across four detection datasets. In instance segmentation tasks on seven core datasets of the BOP challenge, our method outperforms the top RGB methods by 3.6 AP and remains competitive with the best RGB-D method [^1] Code is available at: https://github.com/YoungSean/NIDS-Net.
@@ 1 Introduction
Novel Instance Detection and Segmentation (NIDS) is a crucial task in computer vision aimed at identifying and locating unseen instances in images or videos, given a few examples of each instance. Of particular interest in this work, are applications of NIDS in robotics, e.g., a robot might need to fetch a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the novel object. Ideally, detectors for this task should provide precise bounding boxes and masks of the target instances in a cluttered query image, but standard detection systems, trained on specific classes, often struggle to recognize new objects.
@@
The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the target instance templates, and (3) matching the embeddings of proposals with those of the templates for identification. Recent work [45, 79, 64, 49] has utilized various open-world detectors to obtain object proposals. However, the use of models such as the Segment Anything Model (SAM) [43] or FastSAM [101] often results in the generation of region-based proposals rather than actual object proposals. This misidentification of regions as objects can to lead to object identification issues. For example, a single object may be divided into multiple proposals or some background regions may be mistaken for foreground objects. These false alarms not only slow down the whole detection progress but also spoil the final results. To represent proposals and templates, Shen et al. [79], Nguyen et al. [64], and Lin et al. [49] adopt the $cls$ token of DINOv2  [66, 20] while Li et al. [45] employ a 3D voxel representation. Ideally, for each specific unseen instance, the template embeddings from different views should be similar to each other but markedly different from the embeddings of other instances. Despite these efforts, the challenge persists that the embeddings of different instances remain excessively similar. The development of robust feature embeddings could provide a solution to this issue.
@@
To address the limitations of current methods, we propose a framework for NIDS. Initially, we utilize Grounding DINO [52] with the text prompt “objects” on a cluttered query image to obtain bounding boxes for foreground objects. This method capitalizes on the inherent objectness of Grounding DINO using specialized prompts rather than standard labels. Subsequently, we employ SAM to create masks within these bounding boxes, generating precise object proposals comprising both bounding boxes and masks. For instance embeddings, we first extract average foreground features from the patch embeddings of the DINOv2 ViT backbone. We then apply adapters to enhance instance embeddings by clustering similar instances and distancing different ones. The adapters are trained with instance template embeddings via InfoNCE loss [65, 15]. To refine these embeddings, the CLIP-Adapter [28] introduces a residual vector added to the original embedding, which risks overfitting with a few training examples. This addition can spoil and destabilize the embeddings of background objects, disrupting the entire feature space. To mitigate this issue, we introduce a weight adapter that modifies the original embeddings by applying learned weights. Finally, we employ a straightforward matching approach [60], such as stable matching or using $argmax$, to assign instance labels to these proposals.
@@
Our approach has been validated on four detection datasets and seven segmentation datasets. The framework significantly surpasses existing state-of-the-art methods: it demonstrates substantial increases in average precision (AP), with gains of 22.3, 46.2, 10.3, and 24.0 across these detection datasets. Moreover, in instance segmentation tasks on the seven core datasets of the BOP challenge [82], our method outperforms the best RGB and RGB-D methods.
@@
  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that surpasses existing methods across four detection datasets and seven segmentation datasets of BOP challenge.
  • We leverage the objectness of open-world detector Grounding DINO and region recognition of Segmentation Anything Model (SAM) to obtain object proposals.
@@ 2 Related Work
Pretrained models. Large-scale pretraining has demonstrated utility across various downstream tasks. Works such as [32, 90, 55, 24, 54, 10, 66] emphasize large-scale image representation learning, while [72, 46] focus on vision-language pretraining. Additionally, [3, 4, 95] address depth prediction, and [70, 71, 6, 84] tackle large-scale language modeling. Object detection is addressed by [30, 77, 75, 40, 41, 31, 52], while image segmentation is the focus of [33, 86, 42]. Works such as [93, 48, 14, 74] explore text-to-3D generation, and [80, 1] address text-to-video synthesis. Finally, 6DoF pose estimation and tracking are tackled by [87, 21, 47, 85]. These foundational models offer diverse capabilities, with the challenge lying in effectively leveraging their wealth of knowledge for specific use cases, such as novel instance detection. DINOv2 [66, 20] offers robust features to represent unseen instances.
@@
Open-world object detection (OWOD). OWOD tackles real-world scenarios where object detection algorithms are tasked not only with identifying known and labeled objects but also with effectively handling novel and unknown objects absent during inference. This sets OWOD apart from traditional object detection, which confines itself to detecting only predetermined object classes. A key challenge in OWOD lies in detecting and classifying unknown objects that were not part of the training data. Recent approaches such as OWL-ViT [63], DITO [40], RO-ViT [41], YOLO-World [18], Sambor [31], and Grounding DINO [52] offer diverse solutions to address OWOD challenges. OWL-ViT leverages a standard Vision Transformer [24] architecture with minimal modifications, incorporating contrastive image-text pre-training and end-to-end detection fine-tuning. DITO introduces an open-vocabulary detection approach based on detection-oriented image-text pretraining, replacing the classification architecture with a detector architecture during pretraining to better serve region-level recognition needs. RO-ViT proposes cropping and resizing regions of positional embeddings (PE) at pretraining, aligning better with region-level detection fine-tuning. YOLO-World enhances YOLO [75] with open vocabulary detection capabilities through vision-language modeling and large-scale pretraining. Grounding DINO [52] utilizes features from DINO [10] to ground text queries, while Sambor extends SAM [42] to detect arbitrary objects based on human inputs such as category names or reference expressions. We employ Grounding DINO due to its strong performance.
@@
Few-shot object detection (FSOD). FSOD involves detecting occurrences of a novel class, unseen during training, given only a few template examples. De-ViT [100] introduces a region-propagation mechanism for localization, transforming propagated region masks into bounding boxes using a learnable spatial integral layer. FS-DETR [7], the first few-shot detection transformer, extends DETR [9] to process arbitrary numbers of novel objects concurrently while supporting various examples from each class. Other approaches such as [27] prioritize salient representations through angle distance and magnitude dispersion across few-shot samples,  [92] leverages variational autoencoders for data augmentation, and  [59] learns geometry-aware features with an offline simplex equiangular tight frame classifier. Additionally,  [56] generates specific and representative prototypes for each query image, considering both support and query image features.
@@
Unseen object instance segmentation (UOIS). UOIS is a task designed to segment unseen objects in cluttered scenes. UOIS generates object masks without providing labels. UOIS-Net [89] utilizes a two-stage framework, initially generating rough masks using depth information, and then refining them into precise instance masks with RGB data. Both UCN [88] and MSMFormer [57, 58] segment objects by clustering pixel embeddings. Although these methods can generate object proposals, they primarily focus on tabletop settings. In contrast, our approach is designed to handle RGB query images in various environments.
Adapter. The application of adapters atop large pre-trained models has emerged as a prominent strategy, yielding significant improvements across various tasks. Previous works have leveraged adapters to enhance few-shot image classification tasks [28, 97, 68, 99, 81]. Other works have demonstrated effectiveness in language model fine-tuning [36, 37, 98, 29]. Additional works exhibit the promise of adapters in vision tasks including detection and segmentation [17, 16, 69, 91, 23]. For novel instance detection and segmentation, we propose the Weight Adapter to enhance performance. The adapter assigns weights to the embeddings from frozen backbones.
@@ 3 Method
Novel instance detection and segmentation (NIDS), also referred to as open-world detection and segmentation, is a task to locate and label unseen instances within a query image. We will assume that each of the $N$ target instances is represented by $K$ template images $I_{T}\in\mathbb{R}^{K\times 3\times W\times H}$ and their corresponding segmentation masks. During inference, the output of each query image $I_{Q}\in\mathbb{R}^{3\times W\times H}$ provides these instances’ bounding boxes for detection and instance masks for segmentation.
@@ 3.1 Instance Embedding Generation Stage
In our approach, an instance embedding summarizes the pixels of an instance. The objective of this stage is to generate initial instance embeddings. Each of the $N$ instances has $K$ multi-view template images and their ground truth segmentation masks from which we derive template embeddings $E_{T}\in\mathbb{R}^{N\times K\times C}$, where $C$ is the dimension of the embeddings. Given an image and its corresponding mask, we initially extract patch embeddings using the ViT backbone of DINOv2, and subsequently obtain foreground features as specified by the mask. We then perform average pooling on these features. This process, termed Foreground Feature Averaging (FFA), is proposed by [44] to assess object similarity. We employ FFA to generate all initial instance embeddings.
@@ 3.3 Embedding Refinement via an Adapter
Most instance embeddings are well separated. However, some embeddings from distinct instances may cluster together. For exxample, background object embeddings might be similar to targets. Some target instances may resemble each other. To address this issue, we employ learnable adapters to refine the embeddings. We train these adapters using InfoNCE loss [65, 15], aiming to bring the embeddings of the same instance closer together while separating the embeddings of different instances.
@@
Given an instance embedding $f$ and a $MLP$ (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $f_{c}=\alpha\times MLP(f)+(1-\alpha)\times f$. In this equation, $f_{c}$ represents the adapted embedding, which is a linear combination of the transformed embedding $MLP(f)$ and the original embedding $f$. The residual ratio $\alpha$ is set as 0.6.
@@
We include a ReLU layer before the $sigmoid$ function, ensuring that the sigmoid’s output is constrained to the range $[0.5,1)$. Subsequently, each refined embedding is close to its original counterpart. Additionally, $\beta$ allows for broader adjustments within the feature space. Given that cosine similarity is utilized to compare embeddings in the subsequent phase, our adapter alters the interrelationships among the values within each embedding. Further details of adapters are presented in Fig. 2. Qualitative results can be found in the appendix.
@@ 3.4 Matching Stage
Bonus instance score for segmentation. To improve segmentation performance, we incorporate an additional appearance matching score, $s_{\text{appe}}$, as proposed by SAM-6D [49], into the instance scores. $s_{\text{appe}}$ is used to identify objects that are semantically similar yet differ in appearance. For each proposal, we can identify its most similar template $T_{\text{best}}$ according to template scores. $s_{\text{appe}}$ is derived from the patch embeddings of a proposal image $I$ and $T_{\text{best}}$. It quantifies the similarity between the query image and the template in terms of their local features as follows:
$s_{\text{appe}}=\frac{1}{N_{\text{patch}}^{I}}\sum_{j=1}^{N_{\text{patch}}^{I}}\max_{i=1,\ldots,N_{\text{patch}}^{T_{\text{best}}}}\left(CosineSimilarity(f_{I,j}^{\text{patch}},f_{T_{\text{best}},i}^{\text{patch}})\right)$
@@
After matching, we acquire labeled proposals. Each is defined by $\{b_{i},M_{i},o_{i},s_{i}\}$. Here, $b_{i}$ is the bounding box of an instance. $M_{i}$ is the modal mask, which covers the visible object surface [82]. $o_{i}$ is the instance ID. $s_{i}$ is the confidence score. A confidence score threshold $\delta$ can be used to remove incorrect proposals.
4 Experiments
The experiments are conducted on an RTX A5000 GPU. For each dataset, we train both adapters on template embeddings using InfoNCE loss [65, 15] with temperature 0.05. With two linear layers, Both adapters initially decrease the embedding dimension from $C$ to $C/4$, subsequently restoring it to $C$. The Weight Adapter is trained with Adam optimizer at a learning rate of 1e-3 and a batch size of 1024, while the CLIP-Adapter is trained at a learning rate of 1e-4. Dropout is incorporated in the CLIP-Adapter to reduce overfitting. For additional details, please see the Appendix.
4.1 Detection Datasets
High-resolution dataset Shen et al. [79] design a high-resolution dataset for instance detection. This dataset comprises 100 distinct object instances, each represented by 24 photos taken from multiple viewpoints, with each photo having a resolution of $3072\times 3072$ pixels. These instances are integrated into 14 different indoor scenes for testing, captured in even higher resolution ($6144\times 8192$ pixels). The testing images are further classified as easy or hard based on the level of scene clutter and how much the instances are occluded. Easy tags are assigned when objects are sparsely placed, while hard tags are used for more cluttered setups.
Baselines According to Shen et al. [79], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method. Cut-Paste-Learn generates synthetic training images with 2D-box annotations, putting foreground instances in various sizes and aspect ratios onto different backgrounds. This enables training detectors by viewing each instance as a class. They train five detectors: FasterRCNN [76], RetinaNet [51], CenterNet [102], FCOS [83], and the transformer-based DINO [96]. For the non-learned method, SAM is initially used to generate proposals, followed by employing DINO [11] and DINOv2 [66] to generate features for both proposal and template images, ultimately performing proposal matching and selection.
Synthetic-Real test sets Li et al. [45] employ two benchmarks for evaluation. LineMod-Occlusion (LM-O) [5] includes 8 texture-less objects and 1,514 bounding boxes, primarily challenging due to significant occlusion. The YCB-Video (YCB-V) [8] features 21 objects and 4,125 bounding boxes, with pose variation as the main challenge. Since these datasets have real testing images without reference videos, Li et al. [45] generate synthetic videos using CAD models. We sample 16 synthetic template images per object from these videos.
RoboTools benchmark Li et al. [45] introduce RoboTools, a challenging real-world benchmark for robotics. It features 20 distinct instances, 9,109 annotations, and 24 complex scenarios. Compared to previous benchmarks [5, 8], RoboTools offers greater difficulty due to its cluttered backgrounds and significant pose variations. Additionally, the benchmark’s reference videos are made from real images and incorporate realistic lighting conditions, including shadows. 25 real template images of each instance are sampled from reference videos.
Baselines For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, Li et al. [45] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr. Initially, they generate open-world 2D proposals using their detection module [39]. For matching, different methods are employed to select the proposal with the highest score. In OLNDINO and OLNCLIP, they use robust features from pre-trained backbones [11, 73] and cosine similarity for matching. For OLNCorr, a matching head is designed based on correlation [53]. They also employ the class-level one-shot detectors OS2D [67] and BHRL [94]. To address the limitations of traditional 2D methods with pose variations and occlusions, Li et al. [45] introduced VoxDet which is based on a 3D voxel representation. To train these methods, Li et al. [45] developed the open-world instance detection (OWID-10k) dataset as a synthetic instance detection training set. It features 9,901 objects from ShapeNet [12] and ABO [19], each rendered into a 40-frame, 360-degree video using Blenderproc [22].
Evaluation metrics For detection, we evaluate our method with Average Precision (AP). AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05. This metric is the primary evaluation standard in the COCO Object Detection dataset [50]. Additionally, AP50 and AP75 are variations of this metric, where precision is averaged across all instances at IoU thresholds of 0.5 and 0.75, respectively.
4.2 Segmentation Datasets
BOP challenge We test our method using seven datasets from the BOP challenge [82]: LineMod Occlusion (LM-O) [5], T-LESS [34], TUD-L [35], IC-BIN [25], ITODD [26], HomebrewedDB (HB) [38], and YCB-Video [8]. The objects from these datasets vary widely. We use 42 template rendering images from CNOS [64], which are generated via BlenderProc [22].
Baselines We compare our method with ZeroPose [13], CNOS [64], and SAM-6D [49]. These methods utilize proposals from SAM [43] or FastSAM [101] and use the $cls$ token from DINOv2 as the instance embedding for matching. In addition to adjusting the SAM hyperparameters to generate more proposals, SAM-6D enhances its performance by incorporating an appearance score, $s_{\text{appe}}$. Moreover, SAM-6D employs a Geometric Matching Score, $s_{\text{geo}}$, which considers the shapes and sizes of instances during matching by utilizing depth information.
Evaluation metrics For instance segmentation task, we evaluate our method using the Average Precision (AP) metrics. The AP is computed by averaging the precision scores at various IoU thresholds, from 0.50 to 0.95, increasing by 0.05 at each step.

### v2

@@ Abstract
Novel Instance Detection and Segmentation (NIDS) aims at detecting and segmenting novel object instances given a few examples of each instance. We propose a unified, simple yet effective framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment. Leveraging recent advancements in large vision methods, we utilize Grounding DINO and Segment Anything Model (SAM) to obtain object proposals with accurate bounding boxes and masks. Central to our approach is the generation of high-quality instance embeddings. We utilize foreground feature averages of patch embeddings from the DINOv2 ViT backbone, followed by refinement through a weight adapter mechanism that we introduce. We show experimentally that our weight adapter can adjust the embeddings locally within their feature space and effectively limit overfitting in the few-shot setting. This methodology enables a straightforward matching strategy, resulting in significant performance gains. Our framework surpasses current state-of-the-art methods, demonstrating notable improvements of 22.3, 46.2, 10.3, and 24.0 in average precision (AP) across four detection datasets. In instance segmentation tasks on seven core datasets of the BOP challenge, our method is around 4.5 times faster than the leading published RGB method and surpasses it by 3.6 AP. NIDS-Net is about 5.7 times faster than the top RGB-D method while maintaining competitive performance [^1] Project Page: https://irvlutd.github.io/NIDSNet/.
@@ 1 Introduction
Figure 1: We leverage pre-trained vision models for object proposal generation and feature extraction, and introduce a weight adapter to improve pre-trained feature embeddings for novel object instance detection and segmentation.
@@
Novel Instance Detection and Segmentation (NIDS) is a crucial task in computer vision aimed at identifying and locating unseen instances in images or videos, given a few examples of each instance. It also has wide applications in robotics. For example, suppose a robot needs to grasp a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the object. Ideally, detectors for this task should provide precise bounding boxes and masks of the target instances given a query image. However, standard detection methods trained on specific classes often struggle to recognize new objects.
@@
The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the object templates, and (3) matching the embeddings of proposals with those of the templates for identification. Recent work [49, 81, 66, 53] has utilized various open-world detectors to obtain object proposals. However, the use of models such as the Segment Anything Model (SAM) [47] or FastSAM [103] often results in the generation of region-based proposals rather than actual object proposals. This misidentification of regions as objects can lead to object identification issues. For example, a single object may be divided into multiple proposals or some background regions may be mistaken for foreground objects. These false alarms not only slow down the whole detection progress but also spoil the final results.
To embed proposals and templates, some existing works, e.g., [81, 66, 53], adopt the $cls$ token of DINOv2  [68, 20] while others, e.g., Li et al. [49], employ a 3D voxel representation. Ideally, for each specific unseen instance, the template embeddings from different views should be similar to each other but markedly different from the embeddings of other instances. Despite these efforts, the embeddings of different instances may remain similar, and so the challenges persist.
@@
To address the limitations of the existing methods, we propose a framework for NIDS, as depicted in Fig. 1. Initially, we utilize Grounding DINO [56] with the text prompt “objects” on a cluttered query image to obtain bounding boxes for foreground objects. This method capitalizes on the inherent objectness of Grounding DINO using specialized prompts rather than standard labels. Subsequently, we employ SAM to create masks within these bounding boxes, generating precise object proposals comprising both bounding boxes and masks. For instance embeddings, we first extract average foreground features [48] from the patch embeddings of the DINOv2 ViT backbone. We then apply adapters to enhance instance embeddings by clustering similar instances and distancing different ones.
The adapters are trained using the template images only, where InfoNCE loss [67, 15] is applied to the feature embeddings after the adapters. To refine these embeddings, the CLIP-Adapter [29] introduces a residual vector added to the original embedding, which risks overfitting with a few training examples. This addition can spoil and destabilize the embeddings of background objects, disrupting the entire feature space. To mitigate this issue, we, instead, introduce a weight adapter that modifies the original embeddings by applying learned weights. Finally, we employ a straightforward matching approach [62], such as stable matching or using $argmax$, to assign instance labels to these proposals.
@@
Our approach has been validated on four detection datasets and seven segmentation datasets for NIDS. The framework significantly surpasses existing state-of-the-art methods: it demonstrates substantial increases in average precision (AP), with gains of 22.3, 46.2, 10.3, and 24.0 across these detection datasets. Moreover, in instance segmentation tasks on the seven core datasets of the BOP challenge [84], our method outperforms the best published RGB and RGB-D methods.
@@
  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that surpasses existing published methods across four detection datasets and seven segmentation datasets of BOP challenge.
  • We leverage the objectness of the open-world detector Grounding DINO and region recognition of Segmentation Anything Model (SAM) to obtain object proposals.
@@ 2 Related Work
Pretrained Models. Large-scale pretraining has demonstrated utility across various downstream tasks. Works such as [32, 91, 59, 24, 58, 10, 68] emphasize large-scale image representation learning, while [74, 50] focus on vision-language pretraining. Additionally, [3, 4, 96] address depth prediction, and [72, 73, 6, 87] tackle large-scale language modeling. Object detection is addressed by [79, 77, 43, 44, 31, 56], while image segmentation is the focus of [33, 89, 46]. Works such as [94, 52, 14, 76] explore text-to-3D generation, and [82, 1] address text-to-video synthesis. Finally, 6DoF pose estimation and tracking are tackled by [90, 21, 51, 88]. DINOv2 [68, 20] offers robust features to represent unseen instances. These foundational models offer diverse capabilities, with the challenge lying in effectively leveraging their wealth of knowledge for specific use cases, such as novel instance detection. In our work, we leverage pre-trained vision models for NIDS.
Open-World Object Detection (OWOD). OWOD tackles real-world scenarios where object detection algorithms are tasked not only with identifying known and labeled objects but also with effectively handling novel and unknown objects during inference. This sets OWOD apart from traditional object detection, which confines itself to detecting only predetermined object classes. A key challenge in OWOD lies in detecting and classifying unknown objects that were not part of the training data.
@@
Recent approaches such as OWL-ViT [65], DITO [43], RO-ViT [44], YOLO-World [18], Sambor [31], and Grounding DINO [56] offer diverse solutions to address OWOD challenges. For example, OWL-ViT leverages a standard Vision Transformer [24] architecture with minimal modifications, incorporating contrastive image-text pre-training and end-to-end detection fine-tuning. DITO introduces an open-vocabulary detection approach based on detection-oriented image-text pretraining, replacing the classification architecture with a detector architecture during pretraining to better serve region-level recognition needs. Grounding DINO [56] utilizes features from DINO [10] to ground text queries. We employ Grounding DINO due to its strong performance. Similar to our work, Fang et al. [28] obtain object bounding boxes using the text prompt “a rigid object” with Grounding DINO.
@@
Few-Shot Object Detection (FSOD). FSOD involves detecting occurrences of a novel class, unseen during training, given only a few template examples. For example, De-ViT [102] introduces a region-propagation mechanism for localization, transforming propagated region masks into bounding boxes using a learnable spatial integral layer. FS-DETR [7], the first few-shot detection transformer, extends DETR [9] to process arbitrary numbers of novel objects concurrently while supporting various examples from each class. Other approaches such as [27] prioritize salient representations through angle distance and magnitude dispersion across few-shot samples,  [93] leverages variational autoencoders for data augmentation, and  [61] learns geometry-aware features with an offline simplex equiangular tight frame classifier. Additionally,  [60] generates specific and representative prototypes for each query image, considering both support and query image features. Our method is a FSOD method based on region proposal generation and matching. Compared to previous FSOD methods, it is simple yet effective.
@@
Adapters for Pre-trained Models. The application of adapters atop large pre-trained models has emerged as a prominent strategy, yielding significant improvements across various tasks. Previous works have leveraged adapters to enhance few-shot image classification tasks [29, 99, 70, 101, 83]. Other works have demonstrated effectiveness in language model fine-tuning [37, 39, 100, 30]. Additional works exhibit the promise of adapters in vision tasks including detection and segmentation [17, 16, 71, 92, 23]. For novel instance detection and segmentation, we propose the Weight Adapter to enhance performance. The adapter assigns weights to the embeddings from frozen backbones. SENets [38] employ the Squeeze-and-Excitation (SE) blocks to distribute weights. In contrast, our adapter is an independent model with a structural difference from the SE block. For instance, the SE block lacks a ReLU layer prior to the sigmoid function.
@@ 3 Method
Novel Instance Detection and Segmentation (NIDS), also referred to as open-world detection and segmentation, is a task to locate and label unseen instances within a query image. We will assume that each of the $N$ target instances is represented by $K$ template images $I_{T}\in\mathbb{R}^{K\times 3\times W\times H}$ and their corresponding segmentation masks. During inference, the output of each query image $I_{Q}\in\mathbb{R}^{3\times W\times H}$ provides the bounding boxes for detection and instance masks for segmentation of these instances.
@@ 3.1 Instance Embedding Generation Stage
In our approach, an instance embedding summarizes the pixels of an instance. The objective of this stage is to generate initial instance embeddings. Each of the $N$ instances has $K$ multi-view template images and their ground truth segmentation masks from which we derive template embeddings $E_{T}\in\mathbb{R}^{N\times K\times C}$, where $C$ is the dimension of the embeddings. Given an image and its corresponding mask, we initially extract patch embeddings using the ViT backbone of DINOv2 [68], and subsequently obtain foreground features as specified by the mask. We then perform average pooling on these features. This process, termed Foreground Feature Averaging (FFA), is proposed by [48] to assess object similarity. We employ FFA to generate all initial instance embeddings using the object templates.
@@ 3.3 Embedding Refinement via an Adapter
Most instance embeddings are well separated. However, some embeddings from distinct instances may cluster together. For example, background object embeddings might be similar to targets. Some target instances may resemble each other. To address this issue, we employ learnable adapters to refine the embeddings. We train these adapters using the InfoNCE loss [67, 15], aiming to bring the embeddings of the same instance closer together while separating the embeddings of different instances. This training only uses the template images.
@@
Given an instance embedding $f$ and an $MLP$ (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $f_{c}=\alpha\times MLP(f)+(1-\alpha)\times f$. In this equation, $f_{c}$ represents the adapted embedding, which is a linear combination of the transformed embedding $MLP(f)$ and the original embedding $f$. The residual ratio $\alpha$ is set as 0.6.
@@
We include a ReLU layer before the $sigmoid$ function, ensuring that the sigmoid’s output is constrained to the range $[0.5,1)$. Subsequently, each refined embedding is close to its original counterpart. Additionally, $\beta$ allows for broader adjustments within the feature space. Given that cosine similarity is utilized to compare embeddings in the subsequent phase, our adapter alters the interrelationships among the values within each embedding. Further adapter details are presented in Fig. 3.
@@ 3.4 Matching Stage
Bonus Instance Score for Segmentation. To improve segmentation performance, we incorporate an additional appearance matching score, $s_{\text{appe}}$, as proposed by SAM-6D [53], into the instance scores. The final instance scores are computed as the average of $s_{\text{appe}}$ and the initial instance scores. $s_{\text{appe}}$ is used to identify objects that are semantically similar yet differ in appearance. For each proposal, we can identify its most similar template $T_{\text{best}}$ according to template scores. $s_{\text{appe}}$ is derived from the patch embeddings of a proposal image $I$ and $T_{\text{best}}$. It quantifies the similarity between the query image and the best template in terms of their local features ($f_{I,j}^{\text{patch}}$, $f_{T_{\text{best}},i}^{\text{patch}}$) as follows:
$s_{\text{appe}}=\frac{1}{N_{I}^{\text{patch}}}\sum_{j=1}^{N_{I}^{\text{patch}}}\max_{i=1,\ldots,N_{T_{\text{best}}}^{\text{patch}}}\left(\frac{f_{I,j}^{\text{patch}}\cdot f_{T_{\text{best}},i}^{\text{patch}}}{\|f_{I,j}^{\text{patch}}\|_{2}\|f_{T_{\text{best}},i}^{\text{patch}}\|_{2}}\right),$ (3)
where $N^{\text{patch}}$ is the number of patch embeddings.
@@
After matching, we acquire labeled proposals. Each is defined by $\{b_{i},M_{i},o_{i},s_{i}\}$. Here, $b_{i}$ is the bounding box of an instance $M_{i}$ is the modal mask, which covers the visible object surface [84], $o_{i}$ is the instance ID, and $s_{i}$ is the confidence score. A confidence score threshold $\delta$ can be used to remove incorrect proposals.
@@
4 Experiments
The experiments are conducted on an RTX A5000 GPU. For each dataset, we train both adapters on template embeddings using InfoNCE loss [67, 15] with temperature 0.05. With two linear layers, both adapters initially decrease the embedding dimension from $C$ to $C/4$, subsequently restoring it to $C$. The Weight Adapter is trained with the Adam optimizer [45] at a learning rate of 1e-3 and a batch size of 1024, while the CLIP-Adapter is trained at a learning rate of 1e-4. Dropout [34] is incorporated in the CLIP-Adapter to reduce overfitting. For additional details, please see the appendix.
4.1 Detection Datasets
High-resolution Dataset. Shen et al. [81] design a high-resolution dataset for instance detection. This dataset comprises 100 distinct object instances, each represented by 24 photos taken from multiple viewpoints, with each photo having a resolution of $3072\times 3072$ pixels. These instances are integrated into 14 different indoor scenes for testing, captured in even higher resolution ($6144\times 8192$ pixels). The testing images are further classified as easy or hard based on the level of scene clutter and how much the instances are occluded. Easy tags are assigned when objects are sparsely placed, while hard tags are used for more cluttered setups.
According to Shen et al. [81], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method. Cut-Paste-Learn generates synthetic training images with 2D-box annotations, putting foreground instances in various sizes and aspect ratios onto different backgrounds. This enables training detectors by viewing each instance as a class. They train five detectors: FasterRCNN [78], RetinaNet [55], CenterNet [104], FCOS [85], and the transformer-based DINO [98]. For the non-learned method, SAM is initially used to generate proposals, followed by employing DINO [11] and DINOv2 [68] to generate features for both proposal and template images, ultimately performing proposal matching and selection.
Synthetic-Real Test Sets. Li et al. [49] employ two benchmarks for evaluation. LineMod-Occlusion (LM-O) [5] includes 8 texture-less objects and 1,514 bounding boxes, primarily challenging due to significant occlusion. The YCB-Video (YCB-V) [8] features 21 objects and 4,125 bounding boxes, with pose variation as the main challenge. Since these datasets have real testing images without reference videos, Li et al. [49] generate synthetic videos using CAD models. We sample 16 synthetic template images per object from these videos.
RoboTools Benchmark. Li et al. [49] introduce RoboTools, a challenging real-world benchmark for robotics. It features 20 distinct instances, 9,109 annotations, and 24 complex scenarios. Compared to previous benchmarks [5, 8], RoboTools offers greater difficulty due to its cluttered backgrounds and significant pose variations. Additionally, the benchmark’s reference videos are made from real images and incorporate realistic lighting conditions, including shadows. 25 real template images of each instance are sampled from reference videos.
For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, Li et al. [49] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr. Initially, they generate open-world 2D proposals using their detection module [42]. For matching, different methods are employed to select the proposal with the highest score. In OLNDINO and OLNCLIP, they use robust features from pre-trained backbones [11, 75] and cosine similarity for matching. For OLNCorr, a matching head is designed based on correlation [57]. They also employ the class-level one-shot detectors OS2D [69] and BHRL [95]. To address the limitations of traditional 2D methods with pose variations and occlusions, Li et al. [49] introduced VoxDet which is based on a 3D voxel representation. To train these methods, Li et al. [49] developed the open-world instance detection (OWID-10k) dataset as a synthetic instance detection training set. It features 9,901 objects from ShapeNet [12] and ABO [19], each rendered into a 40-frame, 360-degree video using Blenderproc [22].
Evaluation Metrics. For detection, we evaluate our method with Average Precision (AP). AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05. This metric is the primary evaluation standard in the COCO Object Detection dataset [54]. Additionally, AP50 and AP75 are variations of this metric, where precision is averaged across all instances at IoU thresholds of 0.5 and 0.75, respectively.
4.2 Segmentation Datasets
The BOP Challenge. We test our method using seven datasets from the BOP challenge [84]: LineMod Occlusion (LM-O) [5], T-LESS [35], TUD-L [36], IC-BIN [25], ITODD [26], HomebrewedDB (HB) [40], and YCB-Video [8]. The objects from these datasets vary widely. We use 42 template rendering images from CNOS [66], which are generated via BlenderProc [22].
Baselines. We compare our method with ZeroPose [13], CNOS [66], and SAM-6D [53]. These methods utilize proposals from SAM [47] or FastSAM [103] and use the $cls$ token from DINOv2 as the instance embedding for matching. In addition to adjusting the SAM hyperparameters to generate more proposals, SAM-6D enhances its performance by incorporating an appearance score, $s_{\text{appe}}$. Moreover, SAM-6D employs a Geometric Matching Score, $s_{\text{geo}}$, which considers the shapes and sizes of instances during matching by utilizing depth information.
Evaluation Metrics. For instance segmentation task, we evaluate our method using the Average Precision (AP) metrics. The AP is computed by averaging the precision scores at various IoU thresholds, from 0.50 to 0.95, increasing by 0.05 at each step.